# Line Defects with a Cusp in Fermionic CFTs

### v1

@@ 3.2 Casimir Energy
$\begin{split}\Gamma_{h_{*}\hat{m}_{1},h_{*}\hat{m}_{2}}\sim\frac{1}{\theta}\Bigg[&-\frac{h_{*}^{2}(\hat{m}_{1}\cdot\hat{m}_{2})}{4\pi}-\frac{h_{*}^{2}(\hat{m}_{1}\cdot\hat{m}_{2})}{4\pi}\left(\frac{\epsilon}{2}\log(\pi e^{\gamma})-\frac{g_{1*}^{2}N}{32\pi^{2}}\log(\pi e^{\gamma+2})-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64\pi e^{\gamma-1})\right)\\ &-\left(\epsilon\frac{h_{*}^{2}}{8\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{128\pi^{4}}\right)2\pi(\hat{m}_{1}\cdot\hat{m}_{2})\log 2+\frac{g_{2*}h_{*}^{4}}{96\pi^{3}}(1-\log 2)(\hat{m}_{1}\cdot\hat{m}_{2})\\ &+\frac{g_{2*}h_{*}^{4}}{1024\pi}\left(\frac{1+2(\hat{m}_{1}\cdot\hat{m}_{2})^{2}}{3}\right)\Bigg]\end{split}$ (3.14)
@@
$\begin{split}\mathcal{E}_{h_{*}\hat{m}_{1},h_{*}\hat{m}_{2}}=\Bigg[&-\frac{h_{*}^{2}(\hat{m}_{1}\cdot\hat{m}_{2})}{4\pi}-\frac{h_{*}^{2}(\hat{m}_{1}\cdot\hat{m}_{2})}{4\pi}\left(\frac{\epsilon}{2}\log(\pi e^{\gamma})-\frac{g_{1*}^{2}N}{32\pi^{2}}\log(\pi e^{\gamma+2})-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64\pi e^{\gamma-1})\right)\\ &-\left(\epsilon\frac{h_{*}^{2}}{8\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{128\pi^{4}}\right)2\pi(\hat{m}_{1}\cdot\hat{m}_{2})\log 2+\frac{g_{2*}h_{*}^{4}}{96\pi^{3}}(1-\log 2)(\hat{m}_{1}\cdot\hat{m}_{2})\\ &+\frac{g_{2*}h_{*}^{4}}{1024\pi}\left(\frac{1+2(\hat{m}_{1}\cdot\hat{m}_{2})^{2}}{3}\right)\Bigg]\end{split}$ (3.15)
@@ 3.3 Two point functions of the displacement and tilt operators
$\log\frac{Z_{\hat{m}\hat{m}}(\theta)}{Z_{CFT}}=\log\Bigg[\int[D\mathbf{\phi}D\Psi]e^{-S_{0}}e^{\delta S}\Bigg]$ (3.22)
@@
Figure 3.4: Setting up the coordinate system to compute the two point function of the displacement operator.
@@
$\begin{split}\Gamma^{\prime\prime}(\pi)&=-\frac{h_{*}^{2}}{12\pi^{2}}-\frac{h_{*}^{2}}{12\pi^{2}}\left(\frac{\epsilon}{2}\log(\pi e^{\gamma})-\frac{g_{1*}^{2}N}{32\pi^{2}}\log(\pi e^{\gamma+2})-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64\pi e^{\gamma-1})\right)\\ &-\frac{2}{9}\left(\epsilon\frac{h_{*}^{2}}{8\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{128\pi^{2}}\right)+\frac{g_{2*}h_{*}^{2}}{1152\pi^{8}}(4\pi^{4}-2\pi^{2}-2\pi^{4}\log 8-12\pi^{2}\zeta(3))\\ &+\frac{g_{2*}h_{*}^{4}}{512\pi^{8}}\sigma_{2}\end{split}$ (3.27)
@@
$\begin{split}C_{D}&=\frac{h_{*}^{2}}{2\pi^{2}}+\frac{h_{*}^{2}}{2\pi^{2}}\left(\frac{\epsilon}{2}\log(\pi e^{\gamma})-\frac{g_{1*}^{2}N}{32\pi^{2}}\log(\pi e^{\gamma+2})-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64\pi e^{\gamma-1})\right)\\ &+\frac{4}{3}\left(\epsilon\frac{h_{*}^{2}}{8\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{128\pi^{2}}\right)-\frac{g_{2*}h_{*}^{4}}{192\pi^{8}}(4\pi^{4}-2\pi^{2}-2\pi^{4}\log 8-12\pi^{2}\zeta(3))\\ &+\frac{g_{2*}h_{*}^{4}}{96\pi^{6}}(1+6\zeta(3))\end{split}$ (3.29)
@@
Table 3.2: Numerical values for $C_{D}$ in $d=3$, for some special models
@@
Figure 3.5: Setting up the coordinate system in internal $O(N_{s})$ space to calculate the two point function of the tilt operator.
@@
$\begin{split}h_{*}^{2}\mathcal{N}_{t}^{2}&=\frac{h_{*}^{2}}{4\pi^{2}}\\ &+\frac{h_{*}^{2}}{4\pi^{2}}\left(\frac{\epsilon}{2}\log(\pi e^{\gamma})-\frac{g_{1*}^{2}N}{32\pi^{2}}\log(\pi e^{\gamma+2})-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64\pi e^{\gamma-1})\right)\\ &+\left(\epsilon\frac{h_{*}^{2}}{4\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{64\pi^{2}}\right)\\ &+\frac{g_{2*}h_{*}^{4}}{768\pi^{8}}(-8\pi^{4}+4\pi^{4}\log 8+24\pi^{2}\zeta(3))\\ &-\frac{g_{2*}h_{*}^{4}}{32\pi^{6}}\zeta(3)\end{split}$ (3.38)
@@ A.1 Calculation of the fermion loop diagram
Figure A.1: Fermion Loop diagram
@@
Figure A.2: Scalar line with a fermion loop : Value denoted as ($GNY_{flat}$) from the flat space result and just as (GNY) from the direct cylinder computation

### v2

@@ 3.2 Casimir Energy
$\begin{split}\Gamma_{h_{*}\hat{m}_{1},h_{*}\hat{m}_{2}}\sim\frac{1}{\theta}\Bigg[&-\frac{h_{*}^{2}(\hat{m}_{1}\cdot\hat{m}_{2})}{4\pi}-\frac{h_{*}^{2}(\hat{m}_{1}\cdot\hat{m}_{2})}{4\pi}\left(-\frac{g_{1*}^{2}N}{16\pi^{2}}-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64e^{-1})\right)\\ &-\left(\epsilon\frac{h_{*}^{2}}{8\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{128\pi^{4}}\right)2\pi(\hat{m}_{1}\cdot\hat{m}_{2})\log 2+\frac{g_{2*}h_{*}^{4}}{96\pi^{3}}(1-\log 2)(\hat{m}_{1}\cdot\hat{m}_{2})\\ &+\frac{g_{2*}h_{*}^{4}}{1024\pi}\left(\frac{1+2(\hat{m}_{1}\cdot\hat{m}_{2})^{2}}{3}\right)\Bigg]\end{split}$ (3.13)
@@
$\begin{split}\mathcal{E}_{h_{*}\hat{m}_{1},h_{*}\hat{m}_{2}}=\Bigg[&-\frac{h_{*}^{2}(\hat{m}_{1}\cdot\hat{m}_{2})}{4\pi}-\frac{h_{*}^{2}(\hat{m}_{1}\cdot\hat{m}_{2})}{4\pi}\left(-\frac{g_{1*}^{2}N}{16\pi^{2}}-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64e^{-1})\right)\\ &-\left(\epsilon\frac{h_{*}^{2}}{8\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{128\pi^{4}}\right)2\pi(\hat{m}_{1}\cdot\hat{m}_{2})\log 2+\frac{g_{2*}h_{*}^{4}}{96\pi^{3}}(1-\log 2)(\hat{m}_{1}\cdot\hat{m}_{2})\\ &+\frac{g_{2*}h_{*}^{4}}{1024\pi}\left(\frac{1+2(\hat{m}_{1}\cdot\hat{m}_{2})^{2}}{3}\right)\Bigg]\end{split}$ (3.14)
@@ 3.3 Two point functions of the displacement and tilt operators
$\log\frac{Z_{\hat{m}\hat{m}}(\theta)}{Z_{CFT}}=\log\Bigg[\int[D\mathbf{\phi}D\Psi]e^{-S_{0}}e^{-\delta S}\Bigg]$ (3.21)
@@
Figure 3.4: Setup for the calculation of the two point function of the displacement operator.
@@
$\begin{split}\Gamma^{\prime\prime}(\pi)&=-\frac{h_{*}^{2}}{12\pi^{2}}-\frac{h_{*}^{2}}{12\pi^{2}}\left(-\frac{g_{1*}^{2}N}{16\pi^{2}}-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64e^{-1})\right)\\ &-\frac{2}{9}\left(\epsilon\frac{h_{*}^{2}}{8\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{128\pi^{2}}\right)+\frac{g_{2*}h_{*}^{2}}{1152\pi^{8}}(4\pi^{4}-2\pi^{2}-2\pi^{4}\log 8-12\pi^{2}\zeta(3))\\ &+\frac{g_{2*}h_{*}^{4}}{512\pi^{8}}\sigma_{2}\end{split}$ (3.26)
@@
$\begin{split}C_{D}&=\frac{h_{*}^{2}}{2\pi^{2}}+\frac{h_{*}^{2}}{2\pi^{2}}\left(-\frac{g_{1*}^{2}N}{16\pi^{2}}-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64e^{-1})\right)\\ &+\frac{4}{3}\left(\epsilon\frac{h_{*}^{2}}{8\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{128\pi^{2}}\right)-\frac{g_{2*}h_{*}^{4}}{192\pi^{8}}(4\pi^{4}-2\pi^{2}-2\pi^{4}\log 8-12\pi^{2}\zeta(3))\\ &+\frac{g_{2*}h_{*}^{4}}{96\pi^{6}}(1+6\zeta(3))\end{split}$ (3.28)
@@
Table 3.2: Numerical values for $C_{D}$ in $d=3$, for some special models.
@@
Figure 3.5: Setup for the calculation of the two point function of the tilt operator.
@@
$\begin{split}h_{*}^{2}\mathcal{N}_{t}^{2}&=\frac{h_{*}^{2}}{4\pi^{2}}\\ &+\frac{h_{*}^{2}}{4\pi^{2}}\left(-\frac{g_{1*}^{2}N}{16\pi^{2}}-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64e^{-1})\right)\\ &+\left(\epsilon\frac{h_{*}^{2}}{4\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{64\pi^{2}}\right)\\ &+\frac{g_{2*}h_{*}^{4}}{768\pi^{8}}(-8\pi^{4}+4\pi^{4}\log 8+24\pi^{2}\zeta(3))\\ &-\frac{g_{2*}h_{*}^{4}}{32\pi^{6}}\zeta(3)\end{split}$ (3.37)
@@ A.1 Calculation of the fermion loop diagram
Figure A.1: Fermion loop diagram.
@@
Figure A.2: Scalar propagator with a fermion loop.